The small business Administration offers business loans at 4.95% interest compounded monthly for 10 years. If the owner of a restaurant can afford monthly payments of $1000, what is the maximum amount the owner can borrow?
Answer by Boreal(15235)

Can start with the payment formula or P=(r/k)(PV)/[1-(1+(r/k))^-Nk]
and with r=0.0495 and k=12 and N=10 years,
rewrite the formula as P/[1-(1+(r/k))^-Nk]/(r/k)=PV or present value
This is PV=1000*(1-(1+(0.0495/12)^-120))/(0.0495/12)
=$94498.95 or about $95000.
Round at the end, not intermediate steps.